compute for the force of gravity between two 0.7kg masses whos centers are 15 centimeters apart.
1
Expert's answer
2020-04-20T10:42:16-0400
"F=G\frac{m^2}{r^2}"

"F=6.67\cdot10^{-11}\frac{0.7^2}{0.15^2}=1.45\cdot10^{-9}N"